in a trainagle,R & r are the circumradius and inradius respectively then the H.M of the exradius the triangle is  ???????????

Grade:12

1 Answers

jitender lakhanpal
12 years ago

Dear Zyesha,

 let r1 , r2 , r3 be the three exradii 

r1= Δ/(s-a)   ,r2= Δ/(s-b) , r3= Δ/(s-c)---------------1

where  Δ is the area of triangle   and  s = (a+b+c)/2

the harmonic means of the three exradii are

H.M = 3/(1/r1 + 1/r2 + 1/r3) ---------------2

putting the value of r1,r2,r3 from  1  to 2

 H.M= 3Δ/(s-a+s-b+s-c)  

so H.M = 3Δ/s

and we know that

r = Δ/s

so H.M of exradii = 3 r

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

jitender

 askiitians expert